there were 1500 people at a high school football game. student tickets were $2.00 and adult tickets were $3.50. the total receipts for the game were $3825. how many students bought tickets?
let a = the number of adult tickets purchased
let s = the number of student tickets purchased.
a. write a system of equations that can be used to determine the number of adult and student tickets purchased.
b. determine the number of adults tickets sold and the number of student tickets sold.

Answer by stanbon(75887)

there were 1500 people at a high school football game. student tickets were $2.00 and adult tickets were $3.50. the total receipts for the game were $3825. how many students bought tickets?
let a = the number of adult tickets purchased
let s = the number of student tickets purchased.
a. write a system of equations that can be used to determine the number of adult and student tickets purchased.
a+s=1500
2s+3.5a=3825
-------------------
b. determine the number of adults tickets sold and the number of student tickets sold.
Substitute s=1500-a to solve for a, as follows:
2(1500-a)+3.5a=3825
3000-2a+3.5a=3825
1.5a=825
a=550 (number of student tickets purchased)
s=1500-550=950 (number of adult tickets purchased)
----------
Cheers,
stan H.